OpenStudy (anonymous):

Choose the equation of the line passing through the point (3, -2) and parallel to y = 3x - 6. y = 3x + 7 y = 3x - 11 y = -3x + 11 y = -3x - 12

OpenStudy (anonymous):

Parallel lines have the same slope. The equation you were given is in the form y=mx+b where m = slope. So the slope would be 3. now use point slope form to get the equation. y+2=3(x-3) y+2=3x-9 y=3x-11

OpenStudy (anonymous):

OK I think I understand ; thanks I'll try the next one. But still ask for help